# Participation bias in the estimation of heritability and genetic correlation

### v2

@@ 2.4 Genetic correlation between participation and a phenotype
Intuitively, both genetic and non-genetic correlations between participation and the phenotype can lead to PB. The liability score underlying participation can be thought of as a special phenotype. We define $\rho_{g,PB}=\mathrm{Cov}(\boldsymbol{G}\boldsymbol{\beta},\boldsymbol{G}\boldsymbol{\gamma}^{\prime}|X\geq\Phi^{-1}(1-\alpha))/\sqrt{\mathrm{Var}(\boldsymbol{G}\boldsymbol{\beta})\mathrm{Var}(\boldsymbol{G}\boldsymbol{\gamma}^{\prime}|X\geq\Phi^{-1}(1-\alpha))}$, where $\boldsymbol{\gamma}^{\prime}$ is the per-standardized genetic effect sizes in the sample of participants, which maximizes $\mathrm{Corr}(Y,\boldsymbol{G}\boldsymbol{\gamma}^{\prime}|X\geq\Phi^{-1}(1-\alpha))^{2}$. Note that the $\boldsymbol{\beta}$ inside the expression of $\rho_{g,PB}$ is the per-standardized genetic effect of $X$ in the population, which is not affected by PB. It can be shown that (see Supplementary Note 1.4)
$\rho_{g,PB}\simeq\frac{\rho_{G}-\xi(\alpha)\rho h^{2}_{x}}{\sqrt{1-\xi(\alpha)\rho^{2}}\cdot\sqrt{h^{2}_{x}h^{2}_{y,PB}}}.$ (6)
Some comparisons between $\rho_{g,PB}$ and $\rho_{g}$ are in Figure 1(c). Note that if there is no genetic correlation between the phenotype and participation in the population ($i.e.$ $\rho_{g}=0$), the correlation between genetic components in the sample will be in opposite direction of their phenotypic correlation, a form of collider bias.
2.5 Heritability and genetic correlation estimates with PB adjustment
@@
With the results derived above, we provide a method to correct the effects of PB on the estimates of heritability and genetic correlations. Our method integrates two sources of information: i) the test statistics of participation derived from the IBD-based comparisons (ref.16); and ii) the mean shifts of other phenotypes from population to sample. Notably, although genotypes of non-participants are unavailable, population average of many phenotypes are available from sources such as census data. We start with showing how to adjust the estimate of genetic covariance and heritability here.
@@
We define the mean shift of the standardized phenotype between the selected sample and the population as $\delta=\left[\mathbb{E}(Y|X\geq\Phi^{-1}(1-\alpha))-\mathbb{E}(Y)\right]/\sqrt{\mathrm{Var}(Y|X\geq\Phi^{-1}(1-\alpha))}$. We show that in model (1),
$\delta=\frac{\rho\phi\left[\Phi^{-1}(1-\alpha)\right]}{\alpha\sqrt{1-\xi(\alpha)\rho^{2}}}.$ (7)
@@
Given $\alpha$, $\delta$ is a monotonically increasing function of $\rho$ (Figure S1), the correlation of $X$ and $Y$ and $\delta=0$ if $\rho=0$.
@@
In practice, we estimate the mean phenotypic value in both the sample of participants and another cohort that is representative of the population. The observed mean shift $\widehat{\delta}$ is just difference of the two estimates standardized (divided) by the standard deviation of the phenotypic values of the participants. Notably, while most of the mathematics have been derived for variables standardized in the population, phenotype mean shift is standardized with respect to the sample for convenience of application. For notations, we use $\widehat{\rho}_{g}$, $\widehat{\rho}_{G}$, $\widehat{\varphi}_{g}$, and $\widehat{\varphi}_{G}$ to denote the estimates that are affected but not adjusted for PB, and the corresponding estimates after adjustments are denoted as $\widetilde{\rho}_{g}$, $\widetilde{\rho}_{G}$, $\widetilde{\varphi}_{g}$, and $\widetilde{\varphi}_{G}$. Based on (7), we derive the estimate of the phenotypic correlation between $X$ and $Y$:
$\widehat{\rho}=\frac{\alpha\widehat{\delta}}{\sqrt{\xi(\alpha)\alpha^{2}\widehat{\delta}^{2}+\phi\left[\Phi^{-1}(1-\alpha)\right]^{2}}}.$ (8)
By solving $\rho_{G}$ from (2) of Supplementary Note 1.4, we derive the adjusted genetic covariance of $X$ and $Y$:
@@
where $\widehat{h}^{2}_{x}$ is the heritability of the participation liability score estimated with IBD-based method 16; $\widehat{\rho}_{G}$ is the estimated genetic covariance between $X$ and $Y$ ignoring PB. By solving $h^{2}_{y}$ from (3) and substitute parameters with their estimates, the adjusted heritability estimate of $Y$ is:
$\widetilde{h}^{2}_{y}=\widehat{h}^{2}_{y}(1-\xi(\alpha)\widehat{\rho}^{2})+2\xi(\alpha)\widehat{\rho}\widetilde{\rho}_{G}-\xi(\alpha)^{2}\widehat{\rho}^{2}\widehat{h}^{2}_{x}-\frac{\xi(\alpha)}{1-\xi(\alpha)\widehat{h}^{2}_{x}}\left(\widetilde{\rho}_{G}-\xi(\alpha)\widehat{\rho}\widehat{h}^{2}_{x}\right)^{2},$ (10)
where $\widehat{h}^{2}_{y}$ is the estimated heritability of the phenotype ignoring PB. The adjusted genetic correlation of $X$ and $Y$ is:
@@
Similarly, we solve for $\varphi_{G}$ from the approximation of $\varphi_{G,PB}$ in (18) in and Supplementary Note 1.3, and substitute parameters with their estimates, leading to:
@@
where $\widehat{\varphi}_{G}$ is the estimated genetic covariance of $Y_{1}$ and $Y_{2}$ ignoring PB; $\widehat{\rho}_{G_{1}}$ and $\widehat{\rho}_{G_{2}}$ are the estimated genetic covariance of $X$ and $Y_{1}$ and $Y_{2}$ ignoring PB. The $\widehat{\rho}_{1}$ and $\widehat{\rho}_{2}$ are the estimated phenotypic correlation with participation for the two phenotypes derived with (8). The genetic correlation is then:
@@
where the denominator is derived with (10).
@@
The standard errors are estimated with a block jackknife procedure. In theory, if $\delta$ and $h^{2}_{x}$ are estimated accurately, the adjustment could even lead to a reduced variance, $i.e.$ giving an estimate with both reduced bias and smaller standard error.
In practice, the heritability and genetic correlation are often estimated with marker-based methods such as LD score regression (LDSC). With random sampling, LDSC provides unbiased heritability and genetic covariance estimates 2; 18. In the selected sample, however, not only the SNP effects have changed due to PB, but also the LD structure. We show that the LDSC assumptions are violated in the selected sample, which leads to a negative bias to the estimate of $h^{2}_{y,PB}$ (see Supplementary Note 1.5). Nevertheless, since this bias has an opposite direction relative to the bias resulting from assortative mating 5, the two sources of bias would partially cancel each other in real data applications. Currently, we believe that this bias resulting from the violation of the method’s assumption is negligible, and focus on the gap between $h^{2}_{y}$ and $h^{2}_{y,PB}$.
2.6 Simulations
Our simulation study has two goals. First, we validate our theoretical results on the effect of PB on estimates without adjustment. Second, we show that our adjustment reduces the impact of PB. We generated genetic data for $275,000$ unrelated participants (Methods). The genotypes were simulated from a binomial distribution with block-wise correlation structures. Three scenarios with high genetic correlation ($\rho_{g}=0.75$), low genetic correlation ($\rho_{g}=0.25$), and no genetic correlation ($\rho_{g}=0$) were examined. The non-genetic correlation was fixed at $0.5$. The heritability of participation liability score was fixed at $0.125$, and the participation rate ($\alpha$) was set to $0.055$, which matches the participation rate of the UKBB 7; 16. We used LDSC to estimate heritability and genetic correlation.
For heritability, when the true genetic correlation between participation and the related trait $Y$ was $0.75$, PB led to underestimation biases in heritability estimates for both scenarios with $h^{2}_{y}=0.2$ and $h^{2}_{y}=0.5$. When the genetic correlation was $0.25$ or 00, PB resulted in overestimation bias (Table 1). As for the genetic correlation between two phenotypes other than participation ($\varphi_{g}$), we observed an overestimation bias when $\rho_{g_{1}}=\rho_{g_{2}}=0$. The estimate of $\varphi_{g}$ had an opposite sign when $\rho_{g_{1}}=\rho_{g_{2}}=0.75$ and $\varphi_{g}=0.2$ (Table 2).
Then we applied our adjustment method to the estimates. With adjustments, both the heritability and genetic correlation are very close to the true value in the population, with standard errors comparable to those of the original estimates (Tables 1 and 2).For example, when $h^{2}_{y}=0.2$ and $\rho_{g}=0.75$, the estimated heritability of $Y$ in the sample of participants is 0.156 (SE=0.006). After our adjustment, the estimation is 0.199 (SE=0.010). The estimated genetic correlation between $X$ and $Y$ is 0.503 (SE=0.050) before adjustment, and is 0.749 (SE=0.028) after adjustment. As for the genetic correlation between two other phenotypes, when $\rho_{g_{1}}=\rho_{g_{2}}=0.75$, the estimated of $\varphi_{g}$ is 0.218 (SE=0.029). After adjustment, the estimation is 0.502 (SE=0.027), with 0.5 as the true setting in the population.
Table 1: Theoretical and simulation results for the heritability and genetic correlation estimates. The genetic correlation is measured between the liability score of participation and a related phenotype. Standard deviations of 50 repeated simulations are shown in brackets. The heritability of participation liability score is fixed at 0.125. The non-genetic correlation is fixed at 0.5. $h^{2}_{y,LDSC}$ is the LDSC estimate of heritability of $Y$; and $\rho_{g,LDSC}$ is the LDSC estimate of genetic correlation between $X$ and $Y$.
$h^{2}_{y,LDSC}$ $\rho_{g,LDSC}$
Theoretical Simulated Adjusted True Theoretical Simulated Adjusted True
0.156 0.156 (0.006) 0.199 (0.010) 0.2 0.504 0.503 (0.050) 0.749 (0.028) 0.75
0.230 0.229 (0.009) 0.200 (0.011) 0.2 -0.063 -0.059 (0.067) 0.246 (0.062) 0.25
0.254 0.254 (0.009) 0.201 (0.011) 0.2 -0.273 -0.270 (0.062) -0.001 (0.066) 0
0.465 0.469 (0.012) 0.505 (0.016) 0.5 0.624 0.636 (0.043) 0.758 (0.029) 0.75
0.544 0.549 (0.014) 0.503 (0.016) 0.5 0.084 0.084 (0.055) 0.247 (0.054) 0.25
0.563 0.568 (0.013) 0.505 (0.017) 0.5 -0.141 -0.141 (0.061) -0.003 (0.064) 0
Table 2: Theoretical and simulation results for the estimates of the genetic correlation between two phenotypes other than participation. Standard deviations of 50 repeated simulations are shown in brackets. The heritability of participation liability score is fixed at 0.125. The non-genetic correlation is fixed at 0.5. The heritabilities of the two phenotypes are 0.2 and 0.5. $\varphi_{g,LDSC}$ is the genetic correlation between two phenotypes estimated by LDSC.
Setting $\varphi_{g,LDSC}$
$\rho_{g_{1}},\rho_{g_{2}}$ Theoretical Simulated Adjusted True
0.75 0.218 0.219 (0.029) 0.502 (0.027) 0.5
0.25 0.459 0.459 (0.024) 0.498 (0.027) 0.5
0 0.515 0.515 (0.023) 0.500 (0.029) 0.5
0.75 -0.245 -0.243 (0.053) 0.202 (0.047) 0.2
0.25 0.141 0.143 (0.029) 0.200 (0.035) 0.2
0 0.229 0.232 (0.028) 0.203 (0.037) 0.2
2.7 Real data applications on UKBB phenotypes
@@
We applied our adjustment method to 12 UKBB phenotypes including 4 physical measures: body mass index (BMI), height (HGT), waist circumference (WC), and hip circumference (HC); 3 sociodemographic measures: educational attainment (EA), employment status (ES), and income (INC); and 5 lifestyle measures: current smoking status (SMC), previous smoking status (SMP), alcohol consumption (ALC), walking pace (WKP), and walking time (WKT). Detailed description of the data is summarized in Supplementary Note 1.6. We first estimated heritability and genetic correlations across the phenotypes ignoring PB. We also estimated their genetic correlation with participation based on the genome-wide association studies (GWAS) summary statistics derived from IBD-based information16. Then we calculated the mean shift between the UKBB and the Health Survey for England (HSE) dataset (Methods). We use the HSE dataset as a baseline as it incorporated the weighting to account for nonresponse bias.
@@
Figures 2 and 3 show the effect of PB on the estimates of heritability and genetic correlation of the UKBB phenotypes as functions of mean shifts of the phenotypes. The adjusted estimates based on the observed mean shifts between UKBB and HSE data are indicated. Numerical results of the original (unadjusted) and adjusted estimates are shown in Table 3. With adjustments, the genetic components of BMI, WC, HC, EA, ES, INC, SMC, and WKP are significantly correlated with the genetic components of participation. Specifically, we found the heritability estimates have underestimation bias across those phenotypes that are genetically correlated with participation. Underestimation bias for unadjusted estimates is also observed for the absolute values of genetic correlations. In addition, we found that SMC, which previously lacked significant genetic correlation with participation ($p=0.235$), surpassed the significance threshold of 0.05 with adjustment ($p=0.002$). For ALC and WKT, the unadjusted results showed significant genetic correlation with participation ($p=0.006$ and $0.014$). However, with adjustments, the estimated correlation shrunk and no longer statistically significant ($p=0.081$ and $0.090$). We did not observe estimated correlations that switched signs with adjustment for those phenotypes.
@@
We further analyzed the genetic correlation of each pair of the 12 phenotypes (Figure 4). Compared with the original estimates, we did not observe estimates switching signs. Across phenotypes significantly correlated with participation, we found that, without adjustments, all the genetic correlation estimates had underestimation bias with respect to their absolute value. In addition, the adjusted genetic correlation estimates between WC and EA, HC and EA, and WC and INC, were significantly different from the unadjusted estimates with $p$-values of $0.009$, $0.036$, and $0.048$.
@@
Figure 2: Adjusted heritability estimates of 12 UKBB phenotypes. The red dashed line indicates the original heritability estimates. The blue solid line indicates the heritability estimate adjusted for PB. The purple dashed line marks the mean shift between UKBB and HSE.
@@
Figure 3: Adjusted genetic correlation estimates between participation and the 12 UKBB phenotypes. The red dashed line indicates the original genetic correlation estimates. The blue solid line indicates the genetic correlation estimate adjusted for PB. The purple dashed line marks the mean shift between UKBB and HSE.
@@
Table 3: Adjusted heritability and genetic correlation with participation on 12 physical, sociodemographic, and lifestyle measures in the UKBB. The block jackknife standard errors are shown in brackets. For binary traits, the heritabilities of the liability score are calculated. The significant genetic correlations with participation before and after adjustment are highlighted in boldface. The phenotypes that are significantly correlated with participation after adjustment are highlighted in boldface.
@@
Figure 4: The estimates of genetic correlations across 12 UKBB phenotypes before and after adjustment for PB. The lower left panel: The adjusted genetic correlations. The asterisks highlight the significance level of genetic correlation after adjustment (*: $p<0.05$; **∗ ∗: $p<0.05/78$ (Bonferroni correction)). The upper right panel: The difference between the absolute value of the genetic correlation before and after adjustment. The asterisks highlight the significance level of the difference (*: $p<0.05$).

### v3

@@ 2.4 Genetic correlation between participation and a phenotype
Intuitively, both genetic and non-genetic correlations between participation and the phenotype contribute to the overall correlation, which induces PB when selection occurs based on participation liability score. We define $\rho_{g,PB}=\mathrm{Corr}(G_{x},G^{\prime}_{y}|X>t_{\alpha})$, where $G^{\prime}_{y}=a^{\prime}\cdot G_{x}+G_{w}$ is defined in Section 2.2. We note that $\rho_{g,PB}$ is a special case of $\varphi_{g,PB}$, by regarding $Y_{1}$ as $X$, and $Y_{2}$ as $Y$. Specifically, we have $h^{2}_{y_{1}}=h^{2}_{x}$, $h^{2}_{y_{2}}=h^{2}_{y}$, $\rho_{1}=\rho_{g_{1}}=\rho_{e_{1}}=1$, $\rho_{2}=\rho$, $\varphi_{g}=\rho_{g_{2}}=\rho_{g}$, and $\varphi_{e}=\rho_{e_{2}}=\rho_{e}$. It can be shown that (see Supplementary Note 1.4)
$\rho_{g,PB}=\frac{\rho_{G}-\xi(\alpha)h^{2}_{x}\rho}{\sqrt{(1-\xi(\alpha)\rho^{2})(1-\xi(\alpha)h^{2}_{x})}\cdot\sqrt{h^{2}_{x}h^{2}_{y,PB}}},$ (6)
where the denominator is the squared root of the product of the variance explained by $G_{x}$ and $G^{\prime}_{y}$ in the sample of participants, and the numerator is their covariance. When $X$ and $Y$ are positively correlated (i.e. $\rho>0$), the numerator is smaller than $\rho_{G}$ due to the reduced proportion of $G_{x}$ in $G^{\prime}_{y}$ after selection. Some comparisons between $\rho_{g,PB}$ and $\rho_{g}$ are in Figure 2(c). Note that if $\rho_{g}=0$, $\rho_{g,PB}$ will be in the opposite direction of $\rho$ (or $\rho_{e}$) because of collider bias.
2.5 Estimating heritability and genetic correlation with PB adjustment
@@
With the results derived above, we provide a method to correct the effects of PB on the estimates of heritability and genetic correlations. In addition to the standard estimates computed using the sample, our method utilizes two pieces of information: i) the genetic test statistics for participation derived from the IBD-based comparisons 16; and ii) the mean shifts of other phenotypes from population to sample. Notably, although genotypes of non-participants are unavailable, population average of many phenotypes are available from sources such as census data. We show below how to adjust the estimate of genetic covariance and heritability.
@@
We define the mean shift of a standardized phenotype between the selected sample and the population as $\delta=\left[\mathbb{E}(Y|X>t_{\alpha})-\mathbb{E}(Y)\right]/\sqrt{\mathrm{Var}(Y|X>t_{\alpha})}$. While most mathematics have been derived for variables standardized with respect to the population, here the mean shift is standardized with respect to the sample for the convenience of applications. For model (1),
$\delta=\frac{\rho\phi\left(t_{\alpha}\right)}{\alpha\sqrt{1-\xi(\alpha)\rho^{2}}}.$ (7)
@@
Given $\alpha$, $\delta$ is a monotonically increasing function of $\rho$ (Supplementary Figure 1), and given $\rho$, $\delta$ increases as $\alpha$ decreases.
@@
In practice, we estimate the mean phenotypic value in both the sample of participants and another cohort that is representative of the population. The observed mean shift $\widehat{\delta}$ is just the difference of the two estimates standardized with respect to the sample of participants. Based on (7), we derive the estimate of the phenotypic correlation between $X$ and $Y$:
$\widehat{\rho}=\frac{\alpha\widehat{\delta}}{\sqrt{\xi(\alpha)\alpha^{2}\widehat{\delta}^{2}+\phi\left(t_{\alpha}\right)^{2}}}.$ (8)
[…]
@@
where $\widehat{h}^{2}_{x}$ is the SNP heritability of the participation liability score estimated with IBD-based information 16. The genetic covariance between participation liability score and the phenotype can be derived from (21) of Supplementary Note 1.4:
$\widehat{\rho}_{G}=\widehat{\rho}_{g}\sqrt{(1-\xi(\alpha)\widehat{h}^{2}_{x})\cdot\widehat{h}^{2}_{x}\cdot\widehat{h}^{2}_{y}}.$ (10)
By solving $h^{2}_{y}$ from (3) and substituting parameters with their estimates, the adjusted heritability estimate of $Y$ is:
$\widetilde{h}^{2}_{y}=\widehat{h}^{2}_{y}(1-\xi(\alpha)\widehat{\rho}^{2})+2\xi(\alpha)\widehat{\rho}\widetilde{\rho}_{G}-\xi(\alpha)^{2}\widehat{\rho}^{2}\widehat{h}^{2}_{x}-\frac{\xi(\alpha)}{1-\xi(\alpha)\widehat{h}^{2}_{x}}\left(\widetilde{\rho}_{G}-\xi(\alpha)\widehat{\rho}\widehat{h}^{2}_{x}\right)^{2}.$ (11)
The adjusted genetic correlation of $X$ and $Y$ is:
@@
Similarly, we solve for $\varphi_{G}$ from the approximation of $\varphi_{G,PB}$ in (18) in Supplementary Note 1.3, and substitute parameters with their estimates, leading to:
@@
where $\widehat{\varphi}_{G}=\widehat{\varphi}_{g}\sqrt{\widehat{h}^{2}_{y_{1}}\widehat{h}^{2}_{y_{2}}}$ is the estimated genetic covariance of $Y_{1}$ and $Y_{2}$ ignoring PB; $\widehat{\rho}_{G_{1}}$ and $\widehat{\rho}_{G_{2}}$ are the estimated genetic covariance of $X$ and $Y_{1}$ and $Y_{2}$ ignoring PB, which can be derived from (10). The $\widehat{\rho}_{1}$ and $\widehat{\rho}_{2}$ are the estimated phenotypic correlation with participation for the two phenotypes derived with (8). The adjusted genetic correlation estimate is then:
@@
with the denominator computed with (11).
@@
The standard errors are estimated with a block jackknife procedure. In theory, if $\delta$ and $h^{2}_{x}$ are estimated with small standard errors, the adjustment could even lead to a reduced variance, $i.e.$ giving an estimate with both reduced bias and smaller standard error.
In practice, the heritability and genetic correlation are often estimated with marker-based methods such as LD score regression (LDSC). With random sampling, LDSC provides unbiased heritability and genetic covariance estimates 2; 19, under the assumption that SNPs that are physically far apart are in linkage equilibrium (uncorrelated). In the selected sample, however, not only do the SNP effects change due to PB, but also the LD structure. In particular, SNPs with positive effects on the participation liability score $X$ that are uncorrelated in the population would become negatively correlated in the sample. While this change of LD does not affect (3), (5), and (6), among other things, the LDSC estimate of $h^{2}_{y,PB}$ would be negatively biased when $G_{x}$ and $G_{y}$ are correlated (see Supplementary Note 1.5). Note that this bias is mathematically similar to the positive bias induced by assortative mating 5, only that they are in opposite direction. There are ways to further adjust for this bias if deemed necessary (see Supplementary Note 1.5), but we decided not to do so for the UKBB analyses that follow for two reasons. Firstly, in simulations where the population data are generated under the LDSC assumptions, we found that LDSC estimates of the sample parameters, $e.g.$ $h^{2}_{y,PB}$, are only very slightly biased relative to (3), (5), and (6). Furthermore, the adjusted estimates, $\widetilde{h^{2}_{y}}$, $\widetilde{\rho_{g}}$, and $\widetilde{\varphi_{g}}$, calculated using the LDSC sample estimates are also only very slightly biased relative to the population parameters (see Supplementary Note 1.7 and Supplementary Table 1). Secondly, we examined the odd-chromosome and even-chromosome components of the polygenic score for $X$ and detected neither positive nor negative correlation $p>0.05$ with the UKBB data 16 . Thus, given that the genetic components of $X$ and EA are positively correlated 16, it is plausible that this is a case where some assortative mating effect of $X$ partially cancels out the PB effect on distant LD.
2.6 Analyses of UKBB data
@@
We applied our adjustment method to 12 UKBB phenotypes including 4 physical measures: body mass index (BMI), height (HGT), waist circumference (WC), and hip circumference (HC); 3 sociodemographic measures: EA, employment status (ES), and income (INC); and 5 lifestyle measures: current smoking status (SMC), previous smoking status (SMP), alcohol consumption (ALC), walking pace (WKP), and walking time (WKT). Detailed description of the data is summarized in Supplementary Note 1.6. We first estimated heritability and genetic correlations across the phenotypes ignoring PB. We also estimated their genetic correlation with participation based on the genome-wide association studies (GWAS) summary statistics derived from IBD-based information16. Then we calculated the mean shift between the UKBB and the Health Survey for England (HSE) dataset (Methods). We use the HSE dataset as a baseline as it incorporated the weighting to account for nonresponse bias.
@@
Figures 3 and 4 show the effect of PB on the estimates of heritability and genetic correlation of the UKBB phenotypes as functions of mean shifts of the phenotypes. The adjusted estimates based on the observed mean shifts between UKBB and HSE data are indicated. Numerical results of the original (unadjusted) and adjusted estimates are shown in Table 1. With adjustments, the genetic components of BMI, WC, HC, EA, ES, INC, SMC, and WKP are significantly correlated with the genetic components of participation. Specifically, we found the heritability estimates have underestimation bias across those phenotypes that are genetically correlated with participation. Underestimation bias for unadjusted estimates is also observed for the absolute values of genetic correlations. In addition, we found that SMC, which previously lacked significant genetic correlation with participation ($p=0.235$, two-sided test), surpassed the significance threshold of 0.05 with adjustment ($p=0.002$, two-sided test). For ALC and WKT, the unadjusted results showed significant genetic correlation with participation ($p=0.006$ and $0.014$, two-sided test). However, with adjustments, the estimated correlation shrunk and no longer statistically significant ($p=0.104$ and $0.113$, two-sided test). None of the estimated correlations switched signs with adjustment for those phenotypes.
@@
We further analyzed the genetic correlation of each pair of the 12 phenotypes (Figure 5). No estimate switched signs with adjustment. For phenotypes significantly correlated with participation, absolute values of all the unadjusted genetic correlation estimates appeared to have underestimation bias. In addition, the adjusted genetic correlation estimates between WC and EA, HC and EA, WC and INC, and HC and INC, were significantly different from the unadjusted estimates with $p$-values of $0.006$, $0.021$, $0.028$, and $0.050$ (one-sided test).
@@
Figure 3: Adjusted heritability estimates of 12 UKBB phenotypes. The red dashed line indicates the original heritability estimates. The blue solid line indicates the heritability estimate adjusted for PB across varying mean shifts. The purple dashed line marks the mean shift between UKBB and HSE.
@@
Figure 4: Adjusted genetic correlation estimates between participation and the 12 UKBB phenotypes. The red dashed line indicates the original genetic correlation estimates. The blue solid line indicates the genetic correlation estimate adjusted for PB across varying mean shifts. The purple dashed line marks the mean shift between UKBB and HSE.
@@
Table 1: Adjusted heritability and genetic correlation with participation on 12 physical, sociodemographic, and lifestyle measures in the UKBB. The block jackknife standard errors are shown in brackets. For binary traits, the heritabilities of the liability scores are calculated. The significant genetic correlations with participation before and after adjustment are highlighted in boldface. The phenotypes that are significantly correlated with participation after adjustment are highlighted in boldface.
@@
Figure 5: The estimates of genetic correlations across 12 UKBB phenotypes before and after adjustment for PB. The lower left panel: The adjusted genetic correlations. The asterisks highlight the significance level of genetic correlation after adjustment (*: $p<0.05$; **∗ ∗: $p<0.05/78$ (Bonferroni correction)). Yellow stars highlight cases where significance changed after adjustment, either from significant (threshold of $p<0.05$) to non-significant, or vice versa. The upper right panel: the difference between the absolute value of the genetic correlation before and after adjustment. The size of each circle corresponds to the magnitude of the absolute difference, with larger circles indicating greater differences. The asterisks highlight the significance level of the difference (*: $p<0.05$).